### v1

# Enhancing Urban Visual Place Recognition for Crowdsourced Flood Imagery via LLM-Guided Attention

Abstract
Crowdsourced street-view imagery from social media provides valuable real-time visual evidence of urban flooding and other crisis events, yet it often lacks reliable geographic metadata for emergency response. Existing Visual Place Recognition (VPR) models exhibit substantial performance degradation when applied to such imagery due to visual distortions and domain shifts inherent in cross-source scenarios. This paper presents VPR-AttLLM, a model-agnostic framework that integrates the semantic reasoning and geospatial knowledge of Large Language Models (LLMs) into established VPR pipelines through attention-guided descriptor enhancement. By leveraging LLMs to identify location-informative regions within the city context and suppress transient visual noise, VPR-AttLLM improves retrieval performance without requiring model retraining or additional data. Comprehensive evaluations are conducted on extended benchmarks including SF-XL enriched with real social-media flood images, synthetic flooding scenarios over established query sets and Mapillary photos, and a new HK-URBAN dataset capturing morphologically distinct cityscapes. Integrating VPR-AttLLM with three state-of-the-art VPR models—CosPlace, EigenPlaces, and SALAD—consistently improves recall performance, yielding relative gains typically between 1–3% and reaching up to 8% on the most challenging real flood imagery. Beyond measurable gains in retrieval accuracy, this study establishes a generalizable paradigm for LLM-guided multimodal fusion in visual retrieval systems. By embedding principles from urban perception theory into attention mechanisms, VPR-AttLLM bridges human-like spatial reasoning with modern VPR architectures. Its plug-and-play design, strong cross-source robustness, and interpretability highlight its potential for scalable urban monitoring and rapid geo-localization of crowdsourced crisis imagery.
Keywords
Visual Place Recognition, Geolocalization, Street-View Imagery, Large language Model, GeoAI, Multimodal Fusion
1 Introduction
Street View Imagery (SVI) has become an essential source for urban observation and analysis because it enables detailed, ground-level perspectives on built environments [1, 2]. While platforms such as Google Street View (GSV) provide standardized and extensive spatial coverage, they lack high-frequency temporal updates and cannot capture ephemeral or crisis-related changes in urban scenes—such as flash floods and hurricanes. Consequently, crowdsourced street‑view imagery (SVI) collected through social media and citizen‑reporting platforms has emerged as a valuable complement for timely urban sensing and monitoring [3, 4], enhancing cities’ capacity to assess and respond to rapidly evolving situations, particularly during emergencies [5].
Accurately locating crowdsourced street images is vital for integrating them into spatial decision systems for emergency response and urban resilience planning. This remains challenging because most contributors omit precise geotags for privacy or lack of geospatial literacy for precise locations [6, 7]. Although many cities have deployed official reporting platforms such as New York’s 311 system, persistent reporting friction and limited public awareness constrain participation, leading to biased and incomplete representations of urban issues [8, 9]. During disruptive events such as sudden flooding, citizens often share unstructured photos or videos on social media—data that lack geographic metadata yet offer rich, real‑time visual evidence of urban disruptions (Figure 1). Such imagery complements formal reporting systems by capturing transient, on‑the‑ground scenes that enhance situation awareness. However, current municipal workflows still depend on manual verification, or cross‑referencing by local agencies and journalists, which delays response and limits scalability.
Figure 1: Examples of flood-related street view imagery shared on social media. Top: San Francisco; bottom: Hong Kong. Transient water patterns, reflections, and occlusions substantially alter the street appearance, posing severe challenges for visual place recognition systems trained under normal-weather distributions.
Visual Place Recognition (VPR), also referred to as geo-localization, aims to infer the geographic location of an image by matching it to a geo-tagged reference database. VPR lies at the intersection of computer vision, urban morphology, and environmental perception. Foundational urban perception theories, such as the “Image of the City” [10], "Scale and Psychologies of Space" [11], highlight the role of legibility and structural coherence in spatial recognition, while later computational works have quantified discriminative urban visual elements—demonstrating, for instance, “What makes Paris look like Paris?” [12].
Modern VPR approaches fall broadly into two categories. Retrieval-based methods treat VPR as an image retrieval problem, building city-specific reference databases and retrieving the most similar reference image for a given query—exemplified by NetVLAD [13], CosPlace [14], and more recent SALAD [15]. Classification-based VPR models instead predict a coarse geographic region or city category, allowing global coverage at lower spatial resolution, as demonstrated in PlaNet [16], GeoCLIP [17], and PIGEON [18]. With the progression of deep neural networks, both approaches have achieved significant progress by employing CNN-based backbones (e.g., VGG16, ResNet), vision transformers (e.g., DINO), and multimodal representations (e.g., CLIP).
Despite these advances, robust geo-localization in real-world, crowdsourced SVI remains highly challenging. Even models achieving strong performance on established benchmarks (e.g., SF-XL [14], Pittsburgh-30k [19], Tokyo24/7 [20]) exhibit significant accuracy drops when confronted with distribution shifts, such as extreme weather-induced visual distortions (e.g., flooding, heavy rain) or cross-source variations in data capture devices and environmental conditions. These challenges exemplify a fundamental problem in discriminative AI models: undergeneralization, where model performance degrades considerably when inputs deviate from training distributions [21]. This issue proves particularly acute in VPR for two reasons: (i) extreme weather events generate inherently limited training data due to their rarity, and (ii) the diversity of potential distribution shifts makes scenario-specific retraining computationally prohibitive. Addressing undergeneralization in VPR therefore requires training-free approaches that can dynamically leverage complementary reasoning mechanisms to enhance robustness under challenging conditions without requiring extensive retraining or domain-specific data collection.
The emergence of multimodal Large Language Models (LLMs)—such as GPT-4o, Gemini 2.5, and Qwen-VL—creates new opportunities for resolving these limitations. Beyond text understanding, modern LLMs demonstrate sophisticated scene reasoning and spatial awareness derived from large-scale cross-modal training. They can identify and interpret urban components (e.g., building façades, vegetation, signage, or street furniture) and contextual cues that human observers use to infer place identity. Recent project like GeoIntel [22] show that LLMs can perform coarse geo-reasoning and recognize landmark-rich scenes; however, without integration into established retrieval-based pipelines, such models often lack street-level precision for city-scale VPR tasks.
To address this gap, this paper proposes a VPR enhancement framework with LLM-guided attention (VPR-AttLLM) that leverages LLMs’ semantic reasoning and general geo-knowledge to enhance retrieval-based VPR pipelines under challenging conditions. The central idea is to guide existing pre-trained VPR models with LLM-derived attention map extracted from query images, thereby emphasizing salient regions containing strong location cues (e.g., architectural structure, signage, or horizon geometry) and suppressing noisy regions like flooding street surface (Figure 1). This integration does not require additional training or supporting data, making it model-agnostic and computationally efficient.
Comprehensive experiments demonstrate that the proposed VPR-AttLLM consistently enhances multiple pre-trained VPR models across diverse VPR tests. When applied to CosPlace and EigenPlaces, VPR-AttLLM yields improvements of up to 8% on real flood imagery in the newly constructed HK_Flood dataset and around 4% on the SF_Flood set from San Francisco, while maintaining 1–3% gains under other synthetic flooding and domain-shifted Mapillary scenarios. Notably, the framework demonstrates positive transfer even to the transformer-based SALAD model with a DINOv2 backbone which already achieves strong baseline performance. These results collectively highlight the framework’s ability to reinforce retrieval robustness under severe visual perturbations and cross-source appearance shifts, confirming the generality of LLM-guided attention integration framework.
In summary, this study contributes a novel and generalizable approach for improving VPR robustness under severe appearance or domain shifts by harnessing the semantic reasoning capability and geo-knowledge of large language models. Beyond quantitative gains, the proposed framework highlights the potential for fusing LLM-driven perception with visual retrieval pipelines for resilient urban sensing. The remainder of this paper is organized as follows. Section 2 reviews related work on VPR models and LLM-assisted geo-localization. Section 3 details the proposed VPR-AttLLM framework. Section 4 describes the implementation details, the model performance, and comparative findings. Section 5 discusses implications for multimodal urban sensing, while Section 6 concludes the study.
2 Related Work
2.1 Visual Place Recognition (VPR) Models
Visual Place Recognition (VPR) is commonly formulated as a large-scale image retrieval task, aiming to match a query image to its corresponding geo-referenced counterpart within a database [23, 20]. The primary challenge is to produce compact, discriminative global descriptors that remain robust to changes in viewpoint, illumination, and environmental conditions [24, 25].
Earlier VPR approaches leveraged handcrafted local descriptors (e.g., SIFT, VLAD), but were later superseded by deep convolutional architectures capable of learning aggregation functions end-to-end. A seminal work, NetVLAD, introduced a differentiable VLAD layer on top of CNN backbones to jointly optimize feature representation and place discrimination [13]. Among subsequent CNN-based methods, a major line of progress centered on learnable pooling functions. The Generalized Mean (GeM) pooling extended conventional max- and average-pooling by learning per-channel exponents, thereby enabling more discriminative global descriptors [26]. Building upon this formulation, classification-based VPR frameworks such as CosPlace and EigenPlaces design geo-balanced training strategies that explicitly exploit the strengths of GeM pooling to improve viewpoint and illumination invariance [14, 27]. These works demonstrate how explicit sampling and hierarchical classification amplify the representation capacity of GeM-aggregated descriptors for robust CNN-based VPR.
A parallel direction emphasizes improving aggregation itself through more expressive clustering or token-based formulations. MixVPR introduced feature mixing for better local–global integration [28], while optimal-transport–based architectures such as SALAD extend this concept by learning score-aware cluster aggregation [15]. Transformer-based models, including DINO-ViT and TransVPR, further increasingly adopted with self-attention mechanisms to capture long-range spatial dependencies and offer token-level feature reasoning [29, 30]. Despite these advances, most current methods remain domain-specialized and exhibit performance degradation under severe environmental or cross-source shifts [31].
Beyond architectural innovations in feature extraction and aggregation, complementary post-processing strategies offer performance improvements without requiring model retraining. Among these approaches, Query Expansion (QE) [32] has emerged as a commonly adopted technique in VPR pipelines. Average Query Expansion (AQE) refines the query descriptor by incorporating weighted averages of top-$k$ retrieved database features, with a tunable parameter $\alpha$ controlling the balance between the original query and retrieved candidates. This simple strategy has demonstrated consistent recall improvements with minimal computational overhead across various VPR benchmarks [33, 34]. However, existing post-processing methods have not been systematically evaluated under challenging cross-domain or extreme weather scenarios with severe appearance variations, and semantic-guided query augmentation strategies that leverage high-level scene understanding remain largely unexplored. Addressing these limitations motivates our proposed VPR-AttLLM framework, which introduces language model guided attention modules for both GeM pooling and cluster-based aggregation, thereby unifying these two architectural paradigms under a single interpretive mechanism. By incorporating semantic reasoning into the feature aggregation process, our approach extends beyond conventional post-processing techniques to enable more adaptive query augmentation across diverse environmental conditions.
2.2 Integrating LLMs into Street‑View Perception and Geo‑Localization
In recent years, geo-localization has been widely used to evaluate the geographic knowledge and spatial reasoning ability of LLMs, often through GeoGuessr-like experiments. Several studies benchmark LLMs in inferring countries or cities from given street-view imagery using few-shot prompting [35]. For instance, the GeoIntel agent demonstrates LLM-driven inference of global cities [22]. By leveraging textual or visual signatures, Yin et al. [36] achieved over 60% accuracy within 10 km for disaster image localization during Hurricane Harvey.
Most existing LLM-involved geo-localization frameworks adopt a retrieval-free paradigm, in which models are directly prompted to infer the location name or coordinates from a single image without referencing an explicit database. While this strategy reduces computational cost and avoids large-scale database lookups, it restricts localization performance to coarse resolutions—typically at the city or country level. The high inference cost of querying vast visual databases has thus prevented LLMs from being deployed as retrieval agents capable of precise, street-level matching. As a result, current approaches remain impractical for fine-grained urban sensing, where accurate visual place recognition (VPR) is essential.
In parallel, LLMs have demonstrated strong capabilities in interpreting localized urban scenes and performing multimodal classification tasks. They can infer building ages from street-view imagery [37], estimate inundation depth from flooding images [38], or identify improper on-street waste placement [39] through zero- or few-shot reasoning. However, these applications primarily exploit LLMs as semantic decoders that extract or classify local visual attributes within individual images, rather than as spatially contextual components that influence global feature organization. Additionally, although several relevant experiments have been conducted in LLM’s spatial cognition [40] or geo-knowledge [41], existing tests are still constrained in local scenario inference like demographic indicator prediction. Consequently, while LLMs excel at local semantic interpretation, their potential to guide spatial reasoning or retrieval-level representation in VPR pipelines remains unexplored or underutilized.
Bridging these two gaps—retrieval inefficiency and local-only perception—motivates our proposed VPR-AttLLM framework. Instead of treating LLMs as independent classifiers or scene interpreters, VPR-AttLLM integrates their geographical knowledge and semantic reasoning directly into the global feature aggregation stage of VPR models. By introducing LLM-guided attention modules for both pooling- and cluster-based aggregation, the framework allows LLMs to modulate the weighting of visually informative regions across the entire retrieval procedure. This integration conceptually transforms LLMs from localized predictors into global reasoning agents that enhance cross-condition robustness and fine-resolution geo-localization within existing vision architectures.
3 Methodology
LLMs are capable of inferring the geographic context of a given street view image (SVI) by reasoning about its visual cues and recognizing scene elements that may reveal the underlying location. However, not all image regions contribute equally to this inference. For example, buildings generally carry more discriminative information for localization than roads or skies, and a visually distinctive landmark may provide stronger spatial cues than ordinary residential structures. Conventional VPR models, even when trained on large and diverse datasets, often lack the contextual awareness to assign such importance adaptively—particularly when encountering unseen cities with distinct architectural styles. Moreover, under challenging environmental conditions such as flooding or heavy rain, when query images differ drastically in appearance from the reference database, common VPR models tend to suffer performance degradation due to diminished visual clarity or occlusion. In these cases, an LLM’s capability to reason about spatial salience and refocus on informative regions offers a potential pathway to enhance VPR robustness.
Motivated by this insight, we propose the LLM‑Att framework, an integrated system that leverages an LLM’s broad geo‑semantic knowledge to guide attention in existing pretrained VPR models. The framework consists of two key components: (1) an LLM attention generation module, which guides the LLM to produce spatial attention maps corresponding to salient image regions; and (2) an LLM attention integration module, which fuses external attention maps into various pretrained VPR architectures for inference. Together, these components enable the VPR models to adaptively weight image features according to semantic relevance, improving retrieval performance in challenging queries.
3.1 LLM Attention Generation for SVI Query
In VPR, assigning greater emphasis to informative regions and down-weighting less relevant or noisy areas is crucial for accurately describing a SVI and matching it with images in a reference database. This prioritization aligns with both empirical findings from VPR model training and established principles from urban perception theory. Training dynamics of state-of-the-art VPR models have consistently demonstrated that permanent infrastructure, such as building facades and established vegetation, contributes more reliably to place recognition than transient or generic elements like sky, road surfaces, or moving vehicles. This observation resonates with human spatial cognition patterns documented in environmental psychology literature. However, under challenging visual conditions—such as flooding, extreme illumination shifts, or severe occlusions commonly encountered in crowdsourced social media imagery—the learned feature representations may inadequately capture the optimal region contribution patterns, as models trained on standard street-view datasets have not encountered such distributional shifts during training.
Figure 2: LLM-guided spatial attention weighting for Visual Place Recognition. Axis-based coordinate prompting (top left) guides Gemini 2.5 Flash to generate spatial attention maps with weight values for SF-XL queries. Alternative visual prompting formats are shown for comparison (top right). High weights (1.6–2.0) are assigned to spatially unique landmarks; medium weights (1.0–1.5) to moderately distinctive features. Middle row: single-landmark scenes requiring no weighting. Bottom row: complex urban scenes with weighted attention (shown in orange tags) on distinctive architectural elements (clock towers, ornamental facades) versus generic features (sky, pavement).
Building upon this foundational understanding, urban perception theory provides more granular insights into feature distinctiveness. Lynch’s seminal work [10] on urban legibility identifies landmarks—physical references distinguished by their singularity or memorable form—as critical anchors for spatial recognition, with hierarchical distinctiveness: an architecturally unique civic building possesses greater "landmark value" than a generic residential structure, even though both represent permanent infrastructure. Similarly, Golledge’s [42] wayfinding behavior framework and Sorrows and Hirtle’s [43] landmark salience taxonomy, distinguish between perceptually salient elements that serve as distinctive reference points (e.g., ornate building details, unique signage with high visual or cognitive prominence) and background elements forming the spatial context (e.g., repetitive patterns, standard street furniture with low salience). This suggests that finer-grained differentiation based on spatial uniqueness within the broader urban environment—distinguishing common architectural patterns from rare, localizable features—can further enhance place recognition under adverse conditions.
Our framework operationalizes these principles by prompting the LLM to direct attention toward distinctive urban elements rather than generic scene components based on their spatial uniqueness within the broader cityscape context. Rather than simply segmenting objects, the LLM assesses their contribution to place distinctiveness—effectively distinguishing between common architectural patterns abundant citywide and rare, localizable features. LLMs, with their advanced multimodal reasoning capabilities, offer a novel mechanism to provide such context-aware spatial guidance, particularly when domain shifts degrade the reliability of learned visual representations. The complete prompt is provided in the supplementary material B.
A challenge, however, is that although LLMs can analyze images by extracting features from specific regions, they cannot inherently output explicit spatial attention positions. The task of guiding LLMs to generate coordinates with spatial information within images is known as visual grounding and recent research has explored various strategies for aligning LLM responses to pixel-wise regions within images. For instance, GLaMM [44] introduces a unified framework capable of generating natural language responses seamlessly intertwined with pixel-level segmentation masks, enabling grounded conversation generation. Similarly, VPP-LLaVA [45] leverages visual position prompts to enhance LLaVA’s training performance in visual grounding tasks by explicitly encoding spatial cues.
To operationalize this process, we propose an axis-based visual prompting strategy that positions coordinate axes outside the image frame to provide stable spatial references without altering visual content (Figure 2). This design creates a lightweight and interpretable scaffold that enables the LLM’s text-based reasoning to be spatially grounded during inference. The axis-based approach offers clear advantages over alternative visual prompting methods: grid-based overlays introduce visual artifacts that interfere with the SVIs, while bounding-box references, though providing accurate localization, require additional computation and exhibit failure cases in complex scenes. Consistent with insights from prior multimodal grounding research, including visual position prompting methods [45], our approach benefits from explicit positional cues that strengthen the alignment between textual reasoning and visual localization. Qualitative inspection of the LLM-generated attention maps revealed strong alignment between the highlighted regions and the model’s textual reasoning, indicating consistent spatial grounding. These spatial maps encapsulate intermediate spatial reasoning from the LLM and serve as inputs to the subsequent interpolation process for constructing the full attention map.
Given the LLM-generated dictionary containing $N$ spatial attention points, where each point $i$ is defined by a normalized center coordinate $\mathbf{c}_{i}=(x_{i},y_{i})\in[0,1]^{2}$ and an associated attention weight $w_{i}\in[0,2]$, we construct a smooth attention map $\mathbf{A}\in\mathbb{R}^{H\times W}$ over the feature grid using radial basis function (RBF) interpolation. For each spatial location $(h,w)$ in the feature map, we first compute the normalized grid coordinates $\mathbf{g}_{h,w}=(x_{h,w},y_{h,w})$ where $x_{h,w}=\frac{w}{W-1}$ and $y_{h,w}=\frac{h}{H-1}$. The squared Euclidean distance between grid location $(h,w)$ and each center $i$ is computed as $d^{2}_{h,w,i}=\|\mathbf{g}_{h,w}-\mathbf{c}_{i}\|^{2}$. We then apply a Gaussian RBF kernel with bandwidth parameter $\sigma$ to obtain similarity scores:
$s_{h,w,i}=\exp\left(-\frac{d^{2}_{h,w,i}}{2\sigma^{2}}\right)$ (1)
To construct the final attention map, we compute softmax-normalized interpolation weights and aggregate the attention values from all centers:
$a_{h,w,i}=\frac{s_{h,w,i}}{\sum_{j=1}^{N}s_{h,w,j}},\quad\mathbf{A}_{h,w}=\sum_{i=1}^{N}a_{h,w,i}\cdot w_{i}$ (2)
This formulation ensures that locations closer to specified centers receive attention values closer to their associated weights, while distant regions smoothly blend influences from multiple centers. The bandwidth parameter $\sigma=0.2$ was empirically selected through validation experiments to balance localization precision and smooth spatial transitions. Finally, we clamp the resulting attention values to the valid range $[0,2]$ to prevent extrapolation beyond the specified weight bounds. These refined attention maps are subsequently integrated into multiple VPR backbones to evaluate the overall effectiveness of the proposed framework, providing interpretable and spatially grounded attention representations.
3.2 Attention Integration into VPR models for Inference
The LLM-generated attention maps are integrated into pre-trained VPR models through a general principle: modulate spatial feature contributions in the aggregator according to LLM-derived importance maps while preserving the learned structure of the pre-trained model. This enables adaptive emphasis on semantically discriminative regions without retraining the backbone networks.
3.2.1 General Framework Formulation
Consider a pre-trained VPR model with backbone feature extractor and aggregator module $\mathcal{G}$. For a query image, the backbone produces a spatial feature map $\mathbf{X}\in\mathbb{R}^{C\times H\times W}$, where $C$ denotes the number of feature channels and $H\times W$ represents the spatial resolution. The standard inference pipeline computes the global descriptor as:
$\mathbf{v}=\mathcal{G}(\mathbf{X})\in\mathbb{R}^{D}$ (3)
where $\mathcal{G}$ aggregates spatial features into a $D$-dimensional global descriptor $\mathbf{v}$ for similarity matching against the reference database.
In the VPR-AttLLM framework, we augment this process by incorporating the LLM-generated attention map $\mathbf{A}^{\text{LLM}}\in\mathbb{R}^{H\times W}$ at the final spatial aggregation stage. To control the degree of LLM influence, we introduce a blending coefficient $\alpha\in[0,1]$ that interpolates between the original aggregation and full LLM-guided attention. The attention-modulated inference proceeds as:
$\mathbf{v}^{\text{LLM}}=\mathcal{G}(\mathbf{X};\mathbf{A}^{\text{LLM}},\alpha)$ (4)
where $\mathbf{v}^{\text{LLM}}$ denotes the attention-weighted global descriptor, and $\alpha$ represents LLM impact ratio. When $\alpha=0$, the original pre-trained model behavior is recovered; when $\alpha=1$, full LLM-derived attention is applied. The attention map modulates spatial features within the aggregator $\mathcal{G}$ by scaling each spatial location’s contribution: regions emphasized by the LLM (where $A^{\text{LLM}}_{i,j}>1$) contribute more strongly to the final descriptor, while less discriminative areas (where $A^{\text{LLM}}_{i,j}<1$) are proportionally suppressed. All backbone parameters remain frozen—only the aggregation layer receives attention-modulated inputs during inference.
Figure 3: Overview of the VPR-AttLLM Framework for Enhanced Global Descriptor in Visual Place Recognition. The framework integrates large language model (LLM) attention guidance into existing pre-trained VPR models through a plug-and-play modification of the aggregator module. The pipeline comprises four stages: (1) Query Flooding with LLM Attention Generation: A query street-view image depicting flooding conditions is processed by an LLM (e.g., Gemini-2.5-Flash) with geo-contextual prompts to generate spatial attention weights emphasizing scene elements distinctive for localization in the specified urban context (e.g., San Francisco). (2) Existing Pre-trained VPR Model: Feature extraction using established VPR architectures with diverse backbones (e.g., VGG16, ResNet50, DINOv2) produces a feature map $\mathbf{X}\in\mathbb{R}^{C\times H\times W}$. (3) Blending LLM Attention Map: The aggregator module integrates the LLM attention: (left) original spatial contribution map from the pre-trained model, (center) interpolated LLM attention map ($\sigma=0.2$) highlighting architecturally stable elements, (right) blended contribution map demonstrating enhanced focus on distinctive, localization-relevant structures. (4) Database Matching: The LLM-enhanced query descriptor is matched against unchanged database descriptors from the original VPR model using cosine similarity, enabling seamless integration with existing databases without reprocessing.
Critically, this attention integration applies only to query images at inference time. Reference database images are processed using the original pre-trained model without LLM augmentation. This asymmetric design offers two key advantages: (1) it adjusts the query descriptor formation to emphasize discriminative cues under challenging conditions (cross-source imagery, extreme weather, significant viewpoint variations), and (2) it avoids the computational overhead of processing the entire reference database through the LLM, which would be prohibitively expensive for large-scale deployment.
The specific integration mechanism adapts to the mathematical structure of each aggregator type while maintaining the general principle of attention-based spatial modulation. We demonstrate this across two predominant aggregation paradigms in state-of-the-art VPR architectures: GeM (Generalized Mean) pooling (Section 3.2.2) and cluster-based aggregation (Section 3.2.3). The overall VPR-AttLLM framework is illustrated in Figure 3.
3.2.2 LLM-guided Attention for GeM Pooling
GeM (Generalized Mean Pooling) [26] is widely used as the aggregation function before generating the final descriptor in many VPR models, such as CosPlace, EigenPlaces, CricaVPR. For a feature map $\mathbf{X}\in\mathbb{R}^{C\times H\times W}$, standard GeM pooling computes the output for each channel $c$ as:
$y_{c}=\left(\frac{1}{HW}\sum_{i=1}^{H}\sum_{j=1}^{W}x_{c,i,j}^{p}\right)^{1/p}$ (5)
where $p>1$ is a learnable parameter optimized during training that controls the trade-off between average and max pooling, and $x_{c,i,j}$ denotes the feature value at spatial position $(i,j)$ in channel $c$.
To understand how the pre-trained GeM aggregator implicitly weights different spatial regions, we infer the existing spatial contribution map from the trained parameter $p$. Specifically, for each spatial location $(i,j)$, the effective contribution to the pooled descriptor can be quantified as:
$w^{\text{GeM}}_{i,j}=\left(\frac{1}{HW}\sum_{c=1}^{C}x_{c,i,j}^{p}\right)^{1/p}$ (6)
normalized by its spatial mean to obtain a relative weight map $\mathbf{W}^{\text{GeM}}\in\mathbb{R}^{H\times W}$ that reflects the learned spatial emphasis of the original model.
We then compute a blended attention map that modulates the existing GeM contribution pattern with the LLM attention map $\mathbf{A}^{\text{LLM}}\in\mathbb{R}^{H\times W}$ using blending ratio $\alpha$. The final weight map is defined as:
$\mathbf{W}^{\text{final}}=\mathbf{W}^{\text{GeM}}+\alpha\left(\mathbf{A}^{\text{LLM}}-\mathbf{W}^{\text{GeM}}\right)$ (7)
where this formulation automatically patch region weights where the LLM and pre-trained model disagree: if $\mathbf{A}^{\text{LLM}}_{i,j}\approx\mathbf{W}^{\text{GeM}}_{i,j}$, the adjustment term diminishes, thereby avoiding over-emphasis of already salient regions or excessive suppression of already down-weighted areas.
Utilize the $\mathbf{W}^{\text{final}}$ to modulate spatial aggregation while preserving the original mathematical structure of GeM, the attention-weighted GeM pooling is then computed as:
$y_{c}^{\text{att}}=\left(\frac{\sum_{i=1}^{H}\sum_{j=1}^{W}w^{\text{final}}_{i,j}\cdot x_{c,i,j}^{p}}{\sum_{i=1}^{H}\sum_{j=1}^{W}w^{\text{final}}_{i,j}}\right)^{1/p}$ (8)
Spatial weighting in GeM pooling has been demonstrated valuable for image retrieval tasks [46], and our approach extends this concept by leveraging LLM-derived semantic priors at inference time. This enables seamless integration into existing pre-trained VPR models without retraining the backbone networks, while incorporating the powerful reasoning capabilities of LLMs to address challenging recognition scenarios. Given that GeM pooling is widely adopted in state-of-the-art VPR architectures, the proposed VPR-AttLLM integration framework holds substantial potential for enhancing performance across diverse models and challenging visual place recognition tasks.
3.2.3 LLM-Guided Attention for Cluster Aggregation
While GeM pooling performs global spatial aggregation, another parallel paradigm in VPR employs cluster-based aggregation mechanisms, where spatial features are assigned to multiple semantic clusters via soft probabilities, followed by weighted aggregation to construct descriptors. Notable examples include NetVLAD [13] which employ softmax-based soft assignment, as well as its recent variant SALAD [15], which utilizes optimal transport strategy for more robust cluster assignment. Despite architectural differences, these methods share a common aggregation paradigm: for a feature map $\mathbf{X}\in\mathbb{R}^{C\times H\times W}$ with $n$ spatial units, the model computes cluster features $\mathbf{F}\in\mathbb{R}^{l\times n}$ and an assignment matrix $\mathbf{P}\in\mathbb{R}^{m\times n}$ through learned projections, where $l$ denotes the cluster feature dimension and $m$ represents the number of clusters. The standard aggregation computes each cluster’s representation as:
$\mathbf{V}_{:,k}=\sum_{j=1}^{n}P_{k,j}\cdot\mathbf{F}_{:,j},\quad k=1,\ldots,m$ (9)
where $\mathbf{V}_{:,k}\in\mathbb{R}^{l}$ denotes the aggregated feature vector for cluster $k$, $P_{k,j}$ is the assignment probability of spatial unit $j$ to cluster $k$, and $\mathbf{F}_{:,j}\in\mathbb{R}^{l}$ is the feature vector of spatial unit $j$.
Without modifying pre-trained cluster assignments, our LLM attention map $\mathbf{A}^{\text{LLM}}$ is utilized to modulate feature magnitudes, for controlling spatial unit contributions within their assigned clusters. After reshaping $\mathbf{A}^{\text{LLM}}$ to match the feature map’s spatial structure, denoted as $\mathbf{A}^{\text{LLM}}\in\mathbb{R}^{n}$, with blending ratio $\alpha$, the attention-modulated cluster aggregation then becomes:
$\mathbf{V^{att}}_{:,k}=\sum_{j=1}^{n}P_{k,j}\cdot\left(\alpha\cdot A^{\text{LLM}}_{j}\cdot\mathbf{F}_{:,j}\right),\quad k=1,\ldots,m$ (10)
Here $\mathbf{A}^{\text{LLM}}$ serves as a magnitude modulator that differentially scales feature contributions: distinctive regions identified through multimodal reasoning (e.g., unique architectural landmarks) contribute more strongly to their assigned clusters, while generic or ambiguous regions (e.g., common vegetation or occluding objects) exert proportionally reduced influence. This mechanism enables fine-grained discrimination within the same semantic cluster—for instance, emphasizing a unique historical building over a generic modern structure, even though both belong to the "building" cluster. The resulting attention-modulated cluster features are then processed through the original VPR model’s descriptor construction pipeline to obtain the final LLM-guided descriptor.
4 Experiments
4.1 Datasets
To emulate real-world social media-driven geo-localization scenarios, our evaluation leverages city-wide reference databases that approximate the full urban extent of practical VPR deployment. This configuration enables assessment of LLMs’ broad geographic reasoning capabilities under realistic retrieval conditions.
4.1.1 SF-XL Dataset and Query Construction
Among existing VPR benchmarks (e.g., Pitts250k [20], Nordland, SF-XL [14]), only SF-XL provides urban-scale ground-truth coverage suitable for city-wide evaluation. The SF-XL reference database contains 2.81 million geo-referenced images sampled from 232k unique locations across San Francisco, making it an ideal testbed for assessing VPR performance at metropolitan scale.
Building upon SF-XL, we construct challenging flood-related query sets that prioritize real-world validation while enabling controlled evaluation. Our evaluation strategy consists of three complementary components: (1) Real crowdsourced flooding images: 100 geo-tagged flood photographs manually collected from social media platforms across San Francisco (sf_flood), providing authentic ground truth for extreme weather conditions; (2) Controlled synthetic flooding on benchmark data: Generative augmentation applied to 1,000 original SF-XL queries (sf_v1[F]) to create systematic flooding variations while maintaining exact spatial correspondence with database images; and (3) Contemporary street-view augmentation: 1,000 recent Mapillary images (sf_mapi) with corresponding simulated flood scenes (sf_mapi[F]), bridging temporal gaps in benchmark databases. This multi-faceted approach enables both authentic real-world validation and controlled studies under precisely known correspondences.
Figure 4: Geographic distribution of experimental datasets. (Left) SF-XL dataset covering San Francisco: light gray dots indicate reference database locations ($\sim$2.8M images), light blue shows original benchmark queries (sf_v1, 1000 images), medium blue represents sampled Mapillary images (sf_mapi, 1000 images), and dark blue denotes real flooding images from social media (sf_flood, 100 images). (Right) HK-URBAN dataset covering Hong Kong: reference database contains $\sim$1.15M images from 95,760 locations, with 100 real social media flooding images (hk_flood), 1000 sampled Mapillary images (hk_mapi), and 1000 synthetically generated flooding scenarios (hk_mapi[F]).
To ensure realism while maintaining geometric consistency essential for VPR evaluation, we employ a carefully designed prompting protocol (detailed in Supplementary Material C) using Gemini-2.0-Flash. Following best practices in synthetic data generation for SVI vision tasks [47, 48, 20], our prompts explicitly instruct preservation of original viewpoints and spatial layouts while introducing only appearance degradations: partial inundation, raindrop blur, surface reflections, and atmospheric haze (Figure 5). Critically, all generated images undergo manual verification to confirm that viewpoints and permanent structural elements remain unmodified—only extreme weather effects are added. This view-preserving protocol ensures that any performance differences stem purely from appearance variations rather than geometric inconsistencies, enabling rigorous controlled evaluation of condition-invariant VPR methods.
Figure 5: Examples of simulated flooding applied to original query images, demonstrating controlled degradation while preserving geometric structure and viewpoint.
4.1.2 HK-URBAN: A New Dataset with Distinctive Urban Morphology
To complement SF-XL evaluation and examine performance in a distinct urban morphology, we construct HK-URBAN, a new dataset covering Hong Kong’s primary urban areas with spatial scope comparable to SF-XL. The reference database comprises approximately 1.15 million perspective images extracted from 95,760 geographic locations spanning both Kowloon and Hong Kong Island.
Reference Database Query Dataset Size Description
SF-XL 2.81M images 232K locations sf_flood 100 Real flooding images from social media platforms
sf_v1 1000 Original benchmark queries
sf_v1[F] 1000 Synthetic flooding applied to benchmark queries
sf_mapi 1000 Mapillary images sampled across database coverage
sf_mapi[F] 1000 Synthetic flooding applied to Mapillary samples
HK-URBAN 1.15M images 95.8K locations hk_flood 100 Real flooding images from social media platforms
hk_mapi 1000 Mapillary images sampled across database coverage
hk_mapi[F] 1000 Synthetic flooding applied to Mapillary samples
Table 1: Summary of SF-XL and HK-URBAN datasets used in experiments.
Database construction follows established VPR protocols [20, 14]. Specifically, Google Street View panoramas are sampled at 15 m intervals along all accessible streets within Hong Kong’s central districts. From each panorama, twelve perspective views are rendered with $60^{\circ}$ horizontal field of view (FOV) and $12^{\circ}$ upward pitch to capture building facades and skyline features critical for urban VPR. Query sets mirror those constructed for SF-XL: 100 real social media flood images (hk_flood), 1,000 Mapillary samples (hk_mapi), and 1,000 corresponding simulated flood scenarios (hk_mapi[F]).
HK-URBAN extends evaluation beyond Western urban morphology, enabling assessment of VPR robustness in a dense, high-rise Asian metropolis characterized by distinct architectural styles, vertical urban development, and complex street-level environments. The geographic distribution of all databases and query sets is illustrated in Figure 4, with dataset specifications summarized in Table 1.
4.2 Implementation Details
To demonstrate the generalizability of the proposed VPR-AttLLM framework, we conduct extensive experiments across diverse state-of-the-art VPR architectures with different backbone networks and aggregation mechanisms. For CNN-based paradigms employing Generalized Mean (GeM) pooling as the final aggregator, we evaluate CosPlace [14] and EigenPlaces [27] using their publicly released 512-dimensional descriptors with VGG16 and ResNet backbones, respectively. For Vision Transformer (ViT)-based paradigms with cluster-based aggregation, we evaluate SALAD [15] in both its full configuration (8192D patch features ++ 256D class token) and compact variant (512D patch features ++ 32D class token), where the first component represents spatially aggregated patch-level descriptors and the second captures global scene semantics from the ViT class token.
To contextualize the effectiveness of our training-free, post-processing framework, we compare against Query Expansion (QE), a widely adopted post-retrieval refinement technique in VPR. In our experiments, we implement average QE with weighting parameter $\alpha=0.8$ for the original query descriptor and $(1-\alpha)$ distributed equally among the top-$k$ retrieved reference descriptors, where $k\in\{3,5\}$. This configuration is tested across all three pre-trained models (CosPlace, EigenPlaces, and SALAD) to provide fair comparison against VPR-AttLLM under identical retrieval conditions.
All experiments utilize officially released pre-trained weights without fine-tuning to ensure reproducibility and fair comparison. Backbone networks remain frozen throughout evaluation, with LLM-generated attention maps integrated solely through modified aggregation layers (either LLM-guided GeM pooling or cluster aggregation, as detailed in Section 3.2). Spatial attention maps for all query images are generated using Gemini-2.5-Flash, selected for its extensive geographic knowledge, robust multimodal reasoning capabilities, and particularly strong performance in spatial grounding tasks. The attention map generation method is detailed in Section 3.1.
Following standard evaluation protocols in visual place recognition, performance is measured using Recall@N (with $N=1,5,10$), which denotes the percentage of queries whose top-$N$ retrieved images are within a spatial threshold of the ground-truth location. For the sf_flood and hk_flood query subsets—curated from social media and subject to limited geolocation precision—a 100 m threshold is adopted to accommodate positional noise. For all other queries derived from benchmark or Mapillary sources with accurate geotags, a 25 m threshold is used, aligning with standard practice in VPR evaluations.
4.3 Results
Tables 3 and  2 present comprehensive quantitative results organized around two evaluation contexts: common scenarios representing standard VPR conditions, and flooding scenarios constituting our primary test focus. Common scenarios encompass the benchmark sf_v1 query set alongside two Mapillary-based sets (sf_mapi and hk_mapi), while flooding scenarios comprise synthetic flooding queries (sf_v1[F], sf_mapi[F], hk_mapi[F]) and real social media flood images (sf_flood, hk_flood). All queries are evaluated against the same city-scale reference databases, enabling consistent cross-method and cross-domain comparison.
4.3.1 Performance Under Flooding Scenarios
Under flooding conditions—both real and synthetic—LLM-guided attention demonstrates substantial and consistent improvements over baseline models. Table 2 presents Recall@1, @5, and @10 metrics across five challenging query sets with LLM attention weight $\alpha_{\text{LLM}}\in[0.0,1.0]$, where $\alpha_{\text{LLM}}=0.0$ corresponds to the unenhanced baseline. Figure 6 illustrates Recall@10 variations as $\alpha_{\text{LLM}}$ increases from 0.0 to 1.0 in 0.1 increments, revealing optimal weighting ranges for each scenario.
Results show that VPR-AttLLM consistently improves recall rates across all challenging scenarios, with the most pronounced gains observed in real-world flood imagery (hk_flood, sf_flood) where conventional visual features are most severely degraded. In comparison, Query Expansion (QE)—applied across all test sets for the three VPR backbones—either degrades performance or yields marginal improvements in limited cases, demonstrating no substantial or consistent enhancement relative to VPR-AttLLM.
Model $\alpha_{\textbf{LLM}}$ sf_flood sf_v1[F] sf_mapi[F] hk_flood hk_mapi[F]
@1 @5 @10 @1 @5 @10 @1 @5 @10 @1 @5 @10 @1 @5 @10
VGG16 (512D, CosPlace) 0.0 38.0 44.0 47.0 43.4 53.8 58.3 47.7 59.0 63.3 24.0 37.0 43.0 36.7 49.3 53.2
0.5 41.0 45.0 50.0 44.7 55.0 59.1 49.5 61.0 64.0 23.0 39.0 47.0 37.5 50.1 54.3
1.0 41.0 45.0 49.0 44.8 55.2 59.4 48.3 59.2 63.3 25.0 41.0 50.0 37.1 49.7 53.8
VGG16 (512D, EigenPlaces) 0.0 38.0 40.0 41.0 46.2 53.9 57.4 46.4 57.3 61.0 18.0 33.0 38.0 35.1 49.0 52.9
0.5 37.0 42.0 42.0 47.5 54.4 57.6 47.4 58.6 62.6 18.0 35.0 41.0 35.8 49.9 53.7
1.0 39.0 42.0 44.0 46.7 54.7 58.4 49.1 58.2 62.4 19.0 36.0 41.0 36.4 50.0 53.6
QE(k=3) 38.0 40.0 41.0 45.8 51.9 54.0 46.1 54.8 57.6 18.0 31.0 41.0 34.9 46.3 49.9
QE(k=5) 38.0 40.0 41.0 44.5 53.5 54.9 46.0 55.7 59.0 18.0 32.0 37.0 33.7 48.0 51.7
ResNet50 (512D, CosPlace) 0.0 52.0 57.0 60.0 60.5 68.1 70.9 63.4 74.2 77.3 30.0 45.0 47.0 44.6 58.1 63.4
0.5 53.0 58.0 62.0 61.3 68.5 71.1 64.8 74.8 77.8 30.0 45.0 48.0 45.9 59.7 64.3
1.0 53.0 60.0 64.0 61.1 68.5 70.7 65.5 75.2 78.3 31.0 42.0 48.0 45.6 59.3 63.9
QE(k=3) 50.0 55.0 59.0 60.3 67.1 68.5 63.7 72.3 74.9 33.0 42.0 46.0 44.3 57.1 60.6
QE(k=5) 49.0 56.0 57.0 59.8 67.3 65.8 63.4 73.1 75.4 31.0 44.0 46.0 44.8 57.1 61.6
ResNet (512D, EigenPlaces) 0.0 51.0 60.0 65.0 65.0 71.6 73.4 66.4 75.3 78.5 32.0 43.0 47.0 47.8 60.3 64.7
0.5 52.0 60.0 67.0 65.5 71.6 73.5 67.7 75.7 79.2 31.0 43.0 45.0 48.6 61.4 66.4
1.0 53.0 66.0 67.0 65.7 71.5 73.7 67.3 75.8 78.5 31.0 40.0 48.0 47.3 60.4 65.6
DINOv2 (8192+256D, SALAD) 0.0 63.0 74.0 77.0 64.1 71.1 74.3 71.1 81.3 84.3 60.0 70.0 74.0 66.6 79.7 82.9
0.5 64.0 74.0 77.0 64.1 71.2 74.7 71.6 81.6 84.6 59.0 71.0 75.0 66.7 79.9 83.2
1.0 64.0 75.0 79.0 64.3 71.3 74.6 71.6 81.6 84.8 59.0 72.0 74.0 66.3 79.9 83.3
DINOv2 (512+32D, SALAD) 0.0 50.0 59.0 64.0 51.4 59.5 62.9 54.6 68.4 73.2 42.0 58.0 60.0 53.2 70.3 75.0
0.5 48.0 60.0 63.0 51.7 59.5 63.4 54.7 68.2 72.9 42.0 56.0 61.0 54.6 70.5 74.8
1.0 48.0 60.0 62.0 51.7 59.8 63.4 54.9 68.7 72.8 44.0 52.0 59.0 54.8 70.9 75.5
QE(k=3) 47.0 59.0 59.0 50.6 58.0 60.4 53.7 65.9 70.2 43.0 54.0 60.0 53.4 69.5 73.5
QE(k=5) 48.0 59.0 60.0 50.5 59.0 61.1 54.2 67.6 70.8 42.0 58.0 61.0 53.5 70.2 73.7
Table 2: Recall@N performance on across flooding query sets for six VPR architectures under varying LLM attention weights $\alpha_{\text{LLM}}\in\{0.0,0.5,1.0\}$, where $\alpha_{\text{LLM}}=0.0$ represents the unenhanced baseline. Models include CosPlace and EigenPlaces with VGG16 (512D) and ResNet50 (512D) backbones, and SALAD with DINOv2 backbone in full (8448D) and compact (544D) configurations. Query Expansion (QE) baselines with $k\in\{3,5\}$ are evaluated at $\alpha_{\text{LLM}}=0.0$ for comparison. Recall@1, @5, and @10 metrics are reported across five challenging scenarios. Best performance for each model-query combination is highlighted in bold.
For VGG16-backbone models, performance improvements exceed 3% across real flooding images. Notably, in the hk_flood subset, CosPlace with VGG16 achieves an 8% improvement in Recall@10 at $\alpha_{\text{LLM}}=0.8$, rising from 43.0% to 51.0%. This substantial gain is particularly significant given the model’s complete lack of prior exposure to Hong Kong’s dense high-rise urban morphology, suggesting that LLM-derived semantic priors effectively recover scene structure under severe domain shift and appearance degradation. For less challenging synthetic flooding queries (hk_mapi[F]), steady improvements of 1–2% are observed across wide range of attention weights ($\alpha_{\text{LLM}}\in[0.1,0.9]$).
With ResNet50 as backbone, baseline performance of both CosPlace and EigenPlaces improves by over 10% in absolute recall. VPR-AttLLM maintains consistent gains across both SF-XL and HK-URBAN tests, with Recall@10 improving by up to 5% on sf_flood at $\alpha_{\text{LLM}}=0.9$ and up to 2% on hk_flood. These results further validate the robustness of semantic attention modulation across different backbone configurations and geographic contexts.
For the transformer-based SALAD model with DINOv2 backbone, the pre-trained 8192D++256D descriptor exhibits steady gains of approximately 0.5–1.0% from LLM-guided features, despite its already dense feature encoding from multi-modal DINOv2 pre-training. The compact 512D++32D SALAD variant demonstrates similar performance improvements, though gains are less uniform across all query sets—likely attributable to reduced descriptor capacity for precise cluster assignment. These results indicate that explicit semantic guidance remains valuable even for models with rich cross-modal pre-trained representations, particularly when descriptor dimensionality supports fine-grained spatial encoding.
Figure 6: Recall@10 performance as a function of LLM attention weight $\alpha_{\text{LLM}}$ across six VPR architectures and five flooding query sets. Each subplot shows results for a specific model-backbone combination: CosPlace and EigenPlaces with VGG16 (512D) and ResNet50 (512D) backbones, and SALAD with DINOv2 backbone in both full (8448D) and compact (544D) configurations. The baseline performance ($\alpha_{\text{LLM}}=0.0$) is shown as the leftmost point. VPR-AttLLM consistently improves Recall@10 across nearly all conditions, with real-world flooding scenarios (hk_flood, sf_flood) exhibiting the most substantial gains. Performance remains stable or improving across the full range $\alpha_{\text{LLM}}\in[0.1,1.0]$, with only isolated minor degradation observed at $\alpha_{\text{LLM}}=1.0$ in limited test cases (e.g., hk_flood).
Figure 6 demonstrates that VPR-AttLLM delivers consistent performance improvements across the entire tested range of $\alpha_{\text{LLM}}\in[0.1,1.0]$, with recall rates exceeding baseline performance in nearly all scenarios regardless of attention weight magnitude. Only isolated cases show marginal degradation at the extreme value $\alpha_{\text{LLM}}=1.0$—specifically in hk_flood—suggesting that maximum semantic guidance occasionally over-suppresses fine-grained visual discriminability in certain query-database configurations. Nevertheless, the broad stability of performance gains across diverse attention weights indicates robust semantic-visual integration, with moderate values $\alpha_{\text{LLM}}\in[0.1,0.9]$ typically achieving optimal results while maintaining safe margins from potential degradation zones.
4.3.2 Performance Under Common Scenarios
To assess the generalizability and potential side effects of LLM-guided attention, we evaluate the method on three query sets without appearance degradation: sf_v1, sf_mapi, and hk_mapi. As shown in Table 3, the effects of $\alpha_{\text{LLM}}$ under common scenarios are considerably more subtle than those observed under challenging conditions.
Model $\alpha_{\textbf{LLM}}$ sf_v1 sf_mapi hk_mapi
@1 @5 @10 @1 @5 @10 @1 @5 @10
VGG16 (512D, EigenPlaces) 0.0 70.4 78.7 81.6 72.7 82.4 85.2 58.8 72.9 77.6
0.5 70.2 78.7 81.7 73.6 82.8 85.5 58.3 73.7 78.2
1.0 70.2 78.2 81.1 73.5 82.4 85.0 58.0 73.5 78.0
QE(k=3) 69.9 76.3 78.8 72.6 80.5 82.2 58.0 72.0 75.5
QE(k=5) 70.0 78.6 79.6 73.5 81.6 82.8 58.1 72.2 75.3
ResNet50 (512D, CosPlace) 0.0 78.3 84.6 86.8 82.2 89.8 91.3 67.2 81.7 85.1
0.5 78.0 84.5 87.0 82.5 89.9 91.6 67.1 81.7 85.1
1.0 78.1 84.1 86.9 82.3 90.1 91.6 66.6 81.5 84.6
QE(k=3) 77.4 83.7 85.3 81.9 88.5 90.0 67.1 80.8 83.6
QE(k=5) 77.4 84.2 85.1 82.4 89.2 90.6 67.0 80.9 83.8
DINOv2 (8448D, SALAD) 0.0 88.7 93.5 94.4 85.6 93.2 94.8 76.1 89.4 91.4
0.5 88.8 93.4 94.4 85.0 93.5 94.8 76.1 89.4 91.5
1.0 88.8 93.4 94.5 85.0 93.5 95.1 76.2 89.3 91.5
QE(k=3) 80.7 88.0 88.7 76.0 86.1 88.4 71.1 85.7 87.6
QE(k=5) 80.8 88.9 89.6 75.0 86.7 88.7 71.3 86.1 87.9
Table 3: Recall@N performance on common query sets across six VPR architectures with varying LLM attention weights $\alpha_{\text{LLM}}\in\{0.0,0.5,1.0\}$, where $\alpha_{\text{LLM}}=0.0$ represents the baseline. Query Expansion (QE) results with $k\in\{3,5\}$ top retrievals are included for comparison. Best performance for each model-query combination is shown in bold. VPR-AttLLM maintains stable performance with modest improvements on heterogeneous Mapillary queries (sf_mapi, hk_mapi), while QE consistently degrades baseline performance across all scenarios.
On the benchmark sf_v1 set, where scene composition closely aligns with reference database characteristics and visual patterns are well-captured by baseline descriptors, performance changes remain marginal, with slight degradation occasionally observed at higher $\alpha_{\text{LLM}}$ values. This behavior reflects the limited added value of external semantic guidance when visual features are already discriminative and well-aligned with the reference database.
In contrast, the more heterogeneous Mapillary-based queries (sf_mapi and hk_mapi) reveal modest yet consistent improvements, particularly at moderate attention weights. For both CosPlace and EigenPlaces with VGG16 and ResNet50 backbones, Recall@1 and Recall@5 improve by approximately 1%, indicating that LLM-guided attention provides incremental benefits when camera viewpoints, capture devices, or temporal variations differ from reference conditions. These results demonstrate that while the proposed method does not offer substantial gains under ideal matching conditions, it maintains baseline performance stability while providing small but consistent improvements under subtle domain shifts.
Notably, Query Expansion (QE) fails to improve performance in these common scenarios. Across all tested configurations—both $k=3$ and $k=5$ for top-retrieved candidates—QE consistently degrades baseline performance relative to the original pre-trained descriptors. This degradation stems from QE’s reliance on initial retrieval quality: when top-ranked candidates are geographically distant from the true location, incorporating their features corrupts the query descriptor and compounds retrieval errors. In contrast, VPR-AttLLM’s semantic guidance operates independently of retrieval outcomes, applying attention modulation at the feature aggregation stage before database matching occurs. This design ensures robustness regardless of initial retrieval quality, making the method particularly valuable when query-database domain shifts render conventional similarity metrics unreliable.
Overall, these results validate that LLM-guided attention is most beneficial precisely where it is needed—under challenging visual disruptions—while remaining neutral or mildly helpful in scenarios similar to standard reference databases. The improvements, though moderate in magnitude, are achieved without retraining, additional training data, or architectural modification. Beyond quantitative gains, the explicit attention visualization provides added interpretability, revealing how LLM-driven semantic reasoning reinforces spatial correspondence and perceptual stability in real-world, crowdsourced street-view imagery across diverse urban environments and visual conditions.
4.4 Interpretability Through LLM-Guided Attention
Beyond quantitative improvements, the primary value of the VPR-AttLLM framework lies in its training-free and highly interpretable nature. Without any model fine-tuning or additional optimization, it provides measurable performance gains under extreme or visually unfamiliar conditions, demonstrating strong adaptability to the heterogeneous nature of real-world crowdsourced imagery. This section examines two representative cases—one involving viewpoint shift in common scenarios and another under extreme flooding conditions—to illustrate how LLM-derived semantic guidance complements conventional visual feature extraction.
4.4.1 Case Study 1: Architectural Distinctiveness Under Viewpoint Shift
Figure 7: Qualitative example of VPR-AttLLM in a common scenario with viewpoint shift (sf_mapi). Top row: Query image, with blending LLM attention map’s effects in spatial feature contribution. Medium row: top-5 retrieval results from our VPR-AttLLM enhanced CosPlace, with the correct match highlighted in green. Bottom row: top-5 retrieval results from baseline CosPlace. Despite partial occlusion, the LLM identifies the right building’s distinctive curved bay windows as a strong localization cue (weight 1.6), enabling successful retrieval within top-5 results by redirecting attention away from the fully visible but less architecturally unique left structure.
Figure 7 presents a successful retrieval case from the sf_mapi test set, where the query exhibits moderate viewpoint shift. The original GeM pooling assigns higher feature contributions to the fully visible left building while down-weighting the partially occluded right structure, despite the latter’s distinctive architectural characteristics. This spatial bias arises because conventional pooling mechanisms rely primarily on feature magnitude and spatial extent rather than semantic uniqueness. In contrast, the LLM automatically generates both a weight (1.6) and an interpretable justification for the right building: "Distinctive cream-colored building with curved bay windows and unique entrance, contributing significant architectural uniqueness." By elevating attention to this architecturally unique element along with the adjacent street signage, VPR-AttLLM successfully retrieves the correct reference image within the top-5 results, demonstrating how semantic reasoning can override purely appearance-based feature weighting.
4.4.2 Case Study 2: Stable Feature Selection Under Flooding
Figure 8: Qualitative example of VPR-AttLLM under extreme flooding conditions (sf_flood). Top row: Query image, with blending LLM attention map’s effects in spatial feature contribution. Medium row: top-5 retrieval results from our VPR-AttLLM enhanced EigenPlaces, with the correct match highlighted in green. Bottom row: top-5 retrieval results from baseline EigenPlaces. The baseline model, biased toward the flooded road surface, retrieves visually similar but semantically incorrect tunnel images. VPR-AttLLM identifies the upper marina landscape—boats, docks, and background infrastructure—as offering moderate to high localization cues (weight 1.3), successfully redirecting retrieval toward spatially proximate waterfront references.
Figure 8 illustrates a more challenging case from the sf_flood query set, where severe viewpoint tilt and focal distortion toward the flooded road surface compromise conventional feature extraction. The original GeM pooling fails to capture discriminative information from the upper portion of the image—a marina landscape with boats and docking infrastructure—resulting in retrieval of visually similar but semantically incorrect "gray" tunnel images from the reference database. Due to the scarcity of tilted viewpoint examples in the training data, the baseline model lacks robust priors for such extreme conditions. The LLM, however, readily identifies the upper region as semantically informative, automatically generating the reasoning: "The specific arrangement of boats and visible marina infrastructure (docks, background buildings, poles) offer moderate to high localization cues," and assigns a weight of 1.3 to this region. This semantic reweighting successfully redirects attention toward the marina scene, enabling retrieval of spatially proximate waterfront references and correcting the baseline’s mislocalization.
4.4.3 Geographic Context in Semantic Attention
To examine whether LLM-generated attention maps rely on city-specific semantic knowledge when evaluating architectural distinctiveness, we conduct a controlled ablation study by reversing the city context in attention generation prompts: specifying Hong Kong when processing San Francisco queries and vice versa. This experiment tests whether the same architectural element (e.g., a Victorian facade or a high-rise tower) receives different attention weights depending on its presumed urban context.
Figure 9: Recall@1, @5, and @10 comparison between correct and reversed city context across. Each subplot shows one model evaluated on all flooding query sets. For each metric, bars represent: baseline performance ($\alpha_{\text{LLM}}=0.0$, no hatching), correct city context at $\alpha_{\text{LLM}}=1.0$ (dot hatching) and reversed city context at $\alpha_{\text{LLM}}=1.0$ (diagonal hatching). Reversed prompts yield slightly lower performance than correct context but maintain substantial improvements over baseline under most cases, indicating robustness to imperfect geographic metadata.
Detailed quantitative results across all five flooding scenarios are provided in Supplementary Material A. Figure 9 presents recall results comparison, with reversed-prompt configurations denoted by "(R)" suffix. Reversed city context reduces Recall@10 by around 0.5% compared to correct geographic grounding under most cases. For instance, on sf_mapi[F], CosPlace with ResNet50 achieves 78.3% with correct context versus 77.2% with reversed context ($\Delta=-1.1\%$); EigenPlaces with VGG16 shows 62.4% versus 62.2% ($\Delta=-0.2\%$)).
This modest but consistent degradation reveals two important interpretability insights. First, LLMs do encode city-specific geographic knowledge when assessing landmark distinctiveness. For example, single-family detached houses in Hong Kong are much rarer compared to San Francisco, constituting only a negligible proportion of housing stock limited to luxury enclaves, whereas they represent nearly one-third of San Francisco’s residential units. Conversely, glass curtain-wall towers are more distinctive markers in SF than in HK’s dense skyscraper clusters. This demonstrates that attention maps are not merely generic "interestingness" detectors but rather geographically contextualized evaluations of visual salience. Second, the relatively small magnitude of performance degradation (typically $1.5\%$) and the fact that reversed-prompt configurations still outperform original baselines across most cases indicates the benefit derived from general semantic reasoning about architectural salience—identifying unique facades, landmark structures, and stable visual anchors—beyond precise city-specific knowledge.
These findings carry important practical implications: the robustness of VPR-AttLLM to imperfect city context suggests that even without correct city contexts still provides sufficient general semantic grounding for meaningful improvements, lowering deployment barriers in scenarios where precise metadata may be unavailable or unreliable. More broadly, VPR-AttLLM serves as an effective plug-and-play semantic prior that enhances both spatial reasoning and decision transparency without requiring retraining or domain-specific augmentation. The LLM-derived attention maps provide explicit semantic justifications—emphasizing distinctive urban features such as unique building facades, signage, and infrastructure elements while suppressing visually dominant but semantically uninformative regions (e.g., flooded streets, overexposed sky)—with particular value in challenging conditions where visual features alone prove insufficient. This interpretability through natural language reasoning enables practitioners to audit and refine model behavior in safety-critical applications, bridging the gap between high-performance deep learning and human-understandable decision-making in VPR systems.
5 Discussion
This study demonstrates that training-free semantic guidance from large language models can enhance visual place recognition under challenging conditions, particularly when appearance degradation (e.g., flooding) compounds geographic domain shift. This section synthesizes key findings, examines the framework’s limitations, and identifies opportunities for future integration of multimodal and geospatial reasoning in VPR systems.
5.1 Effectiveness, model agnosticity, and interpretability of LLM-Guided attention
The primary contribution of VPR-AttLLM lies in providing asymmetric query-side enhancement that improves retrieval robustness without reprocessing reference databases—a critical advantage for deployment scenarios involving millions of indexed images. Across five challenging query sets, VPR-AttLLM consistently improved Recall@10 by 1–8%, with the largest gains observed under two conditions: (1) extreme flooding scenarios, where CosPlace achieved up to 8% improvement on real-world flood queries (hk_flood and (2) heterogeneous crowdsourced imagery, where Mapillary queries (sf_mapi[F], hk_mapi[F]) with diverse viewpoints and image quality benefited from 1–2% recall improvements across both cities. Notably, performance gains on Hong Kong Mapillary queries (hk_mapi, hk_mapi[F]) were comparable to those on San Francisco (sf_mapi, sf_mapi[F]), despite models (CosPlace, EigenPlaces) being trained exclusively on San Francisco data. This suggests that LLM-guided attention enhances robustness to query heterogeneity in a cross-city transferable manner. The results validate VPR-AttLLM’s dual capability in handling both extreme appearance degradation (flooding-induced texture distortion) and the intrinsic variability of real-world crowdsourced street-level imagery.
Beyond quantitative improvements, VPR-AttLLM addresses the undergeneralization problem in discriminative models [21], where VPR systems fail under extreme weather conditions such as flooding. The model-agnostic architecture enables seamless integration with diverse VPR backbones without retraining. The late-stage attention fusion strategy operates directly on feature aggregation pahse for pre-trained global descriptors, making it compatible with both CNN-based models (CosPlace, EigenPlaces with VGG16 and ResNet50) and transformer-based architectures (Salad with DINOv2). This plug-and-play property facilitates rapid deployment and experimentation across existing VPR systems, as evidenced by consistent improvements across all three tested architectures. Furthermore, the framework enhances interpretability by providing natural language justifications for attention weights (e.g., "distinctive curved bay windows contributing significant architectural uniqueness"), allowing practitioners to audit which scene elements drive retrieval decisions—a critical requirement for safety-critical geospatial applications such as emergency response or autonomous navigation.
Finally, this work establishes a novel methodological bridge between urban perception theory and computer vision by operationalizing semantic place understanding through LLM-guided attention. The framework’s ability to identify architecturally distinctive elements (e.g., unique facades, landmark structures) and suppress visually dominant but contextually uninformative regions (e.g., front large but common residential building) empirically validates urban theory about place uniqueness in a computational setting. By encoding human-like spatial reasoning—traditionally studied through cognitive mapping and environmental psychology—into feature weighting mechanisms, these interpretable attention maps generated by VPR-AttLLM offer potential value beyond VPR performance: they could serve as data-driven tools for validating urban design theories about place identity and wayfinding [10], or for auditing training dataset biases in visual geolocation systems. By making explicit which scene elements drive model decisions, LLM-guided frameworks bridge the gap between black-box deep learning and human-interpretable spatial reasoning—a critical step toward trustworthy AI in geospatial applications.
5.2 Cross-City Robustness Analysis
VPR models employed in our experiments are pre-trained on geographically constrained datasets: CosPlace and EigenPlaces are trained exclusively on SF-XL, while SALAD is trained on GSV-Cities (predominantly containing urban imagery from the U.S. and Europe). This training-inference geographic disparity provides an opportunity to examine cross-city robustness across both common and challenging flooding scenarios.
[…]
These results suggest two architectural factors influencing cross-city robustness: (1) cluster-based aggregation mechanisms (SALAD) appear more resilient to geographic domain shift than GeM pooling methods, potentially due to their ability to capture diverse visual patterns through learned cluster prototypes; and (2) training on geographically diverse datasets (GSV-Cities) provides moderate improvements over single-city training (SF-XL), though such coverage remains computationally prohibitive at global scale and cannot fully address the long-tail distribution of urban morphologies worldwide.
Notably, VPR-AttLLM maintains consistent improvement margins across both cities despite models’ geographic training biases, indicating that LLM-derived semantic priors offer complementary robustness benefits. This raises a compelling research direction: can LLM guidance be extended from post-hoc attention reweighting to deeper integration within feature extraction stages? Current VPR-AttLLM operates at the aggregation layer, modulating existing local features but unable to recover information absent from the original visual descriptors—an intrinsic limitation of post-processing approaches. Future work integrating semantic reasoning earlier in the training pipeline could potentially achieve more fundamental geographic robustness, moving toward VPR systems with global applicability independent of training city distribution.
5.3 Toward Geography-Aware Multimodal VPR Agents
The broader implication of this work lies in demonstrating that LLMs’ geospatial reasoning—acquired from SVI alone—can meaningfully inform visual recognition tasks in real-world deployment scenarios. This capability opens a path toward geography-aware vision models that leverage LLMs’ implicit knowledge of urban structure (e.g., architectural typologies, street network patterns, landmark hierarchies) to improve cross-domain generalization. For instance, an LLM could dynamically adjust attention priorities based on inferred city context, emphasizing building façades in European historic districts while prioritizing street signage and vertical density cues in Asian megacities. While the current late-fusion strategy treats LLM attention as a post-hoc correction, deeper integration of geospatial reasoning into the learning process holds substantial promise. Future research could explore early-stage integration mechanisms, such as using LLM to guide hard negative mining during contrastive learning, or distilling LLM geospatial priors into lightweight attention modules that can be fine-tuned on target datasets. Such approaches could yield more robust, geography-aware descriptors while maintaining computational feasibility for city-scale deployment.
Beyond visual features, multimodal information present in street-level imagery—particularly text elements such as street signs, storefront names, and directional signage—offers complementary localization cues that remain underutilized in current VPR systems. Although such text regions receive high attention weights in LLM-generated attention maps (Figure 7), they cannot be effectively exploited without corresponding geospatial-aware reference databases. Early works in OCR-augmented VPR [49] and recent adavances in LLM-based multimodal reasoning for image geolocalization [36] suggest that fusing visual global descriptors with text embeddings may yield substantial improvements, particularly in urban environments where signage is prevalent and culturally informative.
Looking forward, LLMs hold potential to serve as orchestrating agents that coordinate multiple retrieval modalities—integrating dense street-level imagery with sparse but semantically rich point-of-interest (POI) data, leveraging tool-augmented reasoning for geospatial database queries, and synthesizing evidence across visual, textual, and structural cues. Such agent-based architectures could address challenging real-world localization tasks in emergency response scenarios (e.g., disaster damage assessment from social media imagery) or autonomous navigation in rapidly changing urban environments, where robust place recognition under extreme conditions is not merely an academic exercise but a critical operational requirement. By bridging human-like spatial reasoning with scalable visual retrieval, this research direction promises to transform VPR from a pattern-matching problem into a truly geography-aware, context-sensitive spatial intelligence capability.
6 Conclusion
This work introduces VPR-AttLLM, a training-free framework that leverages LLM’ geospatial reasoning to enhance visual place recognition under challenging real-world conditions. By generating semantically informed attention maps that emphasize architecturally distinctive features while suppressing visually dominant but contextually uninformative regions, the framework achieves consistent performance improvements (1–8%) across multiple VPR architectures, with the most substantial gains observed in flood-degraded imagery scenarios. Critically, VPR-AttLLM operates as a plug-and-play query-side enhancement that requires no model retraining or reference database reprocessing, enabling immediate deployment on existing city-scale VPR systems while providing interpretable natural language justifications for retrieval decisions.
Beyond immediate technical contributions, this research establishes a methodological pathway for operationalizing urban perception theory within computer vision systems and demonstrates the viability of geography-aware multimodal reasoning for robust place recognition. The framework’s cross-domain effectiveness—particularly on unseen cities and extreme appearance conditions—validates that LLM-derived semantic priors encode transferable spatial intelligence that complements learned visual representations. As street-level imagery continues to proliferate through crowdsourced platforms and autonomous systems, the integration of semantic reasoning with visual retrieval represents a foundational step toward next-generation VPR agents capable of human-aligned, context-sensitive localization in increasingly complex and dynamic built environments. By demonstrating that such capabilities can emerge from pre-trained language models without task-specific fine-tuning, this work opens scalable avenues for interpretable and adaptive approaches to visual geolocalization.
Declaration of generative AI and AI-assisted technologies in the writing process
During the preparation of this work the author(s) used Claude 4.5 (Sonnet) in order to improve language clarity and readability through grammar refinement and word choice suggestions. After using this tool/service, the author(s) reviewed and edited the content as needed and take(s) full responsibility for the content of the published article.
Acknowledgements
The authors gratefully acknowledge partial financial support from the Seed Fund for PI Research – Basic Research (Grant No. 2402101354), University Research Committee, The University of Hong Kong.
Appendix
Appendix A Experiment results for reversed city used in prompt
Model $\alpha_{\textbf{LLM}}$ sf_flood sf_v1[F] sf_mapi[F] hk_flood hk_mapi[F]
@1 @5 @10 @1 @5 @10 @1 @5 @10 @1 @5 @10 @1 @5 @10
VGG16 (512D, EigenPlaces) 0.0 38.0 40.0 41.0 46.2 53.9 57.4 46.4 57.3 61.0 18.0 33.0 38.0 35.1 49.0 52.9
0.5 37.0 42.0 42.0 47.5 54.4 57.6 47.4 58.6 62.6 18.0 35.0 41.0 35.8 49.9 53.7
1.0 39.0 42.0 44.0 46.7 54.7 58.4 49.1 58.2 62.4 19.0 36.0 41.0 36.4 50.0 53.6
0.5 (R) 39.0 41.0 42.0 47.0 54.4 57.7 47.1 59.2 62.3 18.0 32.0 41.0 35.4 49.2 53.8
1.0 (R) 36.0 41.0 42.0 46.5 54.2 57.7 48.6 58.2 62.2 20.0 37.0 42.0 34.5 47.6 51.7
ResNet50 (512D, CosPlace) 0.0 52.0 57.0 60.0 60.5 68.1 70.9 63.4 74.2 77.3 30.0 45.0 47.0 44.6 58.1 63.4
0.5 53.0 58.0 62.0 61.3 68.5 71.1 64.8 74.8 77.8 30.0 45.0 48.0 45.9 59.7 64.3
1.0 53.0 60.0 64.0 61.1 68.5 70.7 65.5 75.2 78.3 31.0 42.0 48.0 45.6 59.3 63.9
0.5 (R) 52.0 59.0 62.0 61.0 68.5 70.9 65.4 74.5 78.0 34.0 46.0 48.0 45.3 59.9 64.5
1.0 (R) 53.0 59.0 64.0 61.1 68.2 70.8 65.5 74.2 77.2 35.0 44.0 51.0 45.7 58.3 63.1
DINOv2 (544D, SALAD) 0.0 50.0 59.0 64.0 51.4 59.5 62.9 54.6 68.4 73.2 42.0 58.0 60.0 53.2 70.3 75.0
0.5 48.0 60.0 63.0 51.7 59.5 63.4 54.7 68.2 72.9 42.0 56.0 61.0 54.6 70.5 74.8
1.0 48.0 60.0 62.0 51.7 59.8 63.4 54.9 68.7 72.8 44.0 52.0 59.0 54.8 70.9 75.5
0.5 (R) 49.0 60.0 63.0 51.2 59.6 63.4 54.9 68.8 73.0 42.0 57.0 63.0 54.2 70.5 74.7
1.0 (R) 48.0 59.0 64.0 51.5 59.7 63.3 54.3 69.9 73.3 41.0 56.0 65.0 54.2 70.2 74.3
Complete ablation study of reversed city context across all challenging query sets and LLM attention weights. "Correct" denotes attention maps generated with accurate geographic context (SF for San Francisco queries, HK for Hong Kong queries); "Reversed" uses inverted city context. Results demonstrate consistent 0.5–1.5% degradation with reversed context while maintaining performance above baseline ($\alpha_{\text{LLM}}=0.0$) across all configurations, validating robustness to imperfect geographic metadata.
Appendix B Prompts used for generating LLM attention
* VPR Geo-localization: Scene Composition Weighting
**TASK:** Your goal is to determine if an image requires spatial weighting for Visual Place Recognition (VPR) and, if so, to provide those weights. Follow this two-step decision process strictly.
STEP 1: DECIDE THE SCENE TYPE
First, classify the image into one of two categories. This is the most critical step.
A) ’Single Landmark’ Image: The image is dominated by one primary, cohesive structure. Criteria: This includes focused shots of a single building’s facade, a prominent monument, a statue, or a bridge, etc. The object acts as a single, unambiguous identifier. Crucial Rule: Even if this single landmark has unique details (e.g., intricate windows, a brand name, specific textures), it must be treated as one whole entity. Do not segment a single building into parts. Applying different weights to parts of one building is incorrect and adds noise. ACTION: If the image is a ’Single Landmark’, STOP. Your entire and only output must be the single string: "None"
B) ’Complex Scene’ Image: The image contains multiple, distinct, and geographically significant objects that contribute differently to localization. Criteria: This includes street corners with several buildings, plazas with fountains and surrounding structures, or street views with varied storefronts and unique street furniture. The combination of these separate objects defines the location. ACTION: If the image is a ’Complex Scene’, PROCEED to Step 2 to generate weights.
STEP 2: GENERATE WEIGHTS (FOR ’COMPLEX SCENE’ IMAGES ONLY) If and only if you identified the image as a ’Complex Scene’ in Step 1, identify 3-8 key regions and assign weights based on their uniqueness for localization in the city. The city is city, and the spatial uniqueness is the uniqueness of the region in the whole city.
**COORDINATE SYSTEM:** Image coordinates range from (0.0, 0.0) at top-left to (1.0, 1.0) at bottom-right Center coordinates represent the focal point of each distinctive region
**WEIGHTING LOGIC:** **HIGH WEIGHTS (1.6-2.0):** Spatially unique in city - Rare architectural details, unusual building features - Distinctive facades, unique structural elements - Sharp, detailed building textures/patterns - Uncommon urban installations or specialized infrastructure
**MEDIUM WEIGHTS (1.0-1.5):** Moderately distinctive in city - Well-detailed standard buildings with clear architectural features - Street-level infrastructure with visible specificity - Clear signage, distinctive storefronts
**LOW WEIGHTS (0.3-0.9):** Common elements in city - Typical building styles (abundant citywide) - Generic street infrastructure, standard traffic elements
**MINIMAL WEIGHTS (0.0-0.2):** Non-localizable in city - Sky, vegetation, generic pavement - People, vehicles, temporary objects - Blurred or indistinct regions
**SELECTION STRATEGY:** Focus on 3-8 most important regions that deviate significantly from default weight (1.0) Prioritize regions with highest uniqueness (>1.5) and lowest uniqueness (<0.5) Skip regions that would receive approximately default weight (0.9-1.1)
**OUTPUT FORMAT:**
        "center": [x_coord, y_coord],
        "weight": weight_value,
        "reasoning": "brief_description"
Appendix C Prompts used for simulating flooding queries
Simulate a realistic slight inundation at this location. Strictly preserve all original buildings and infrastructure. Maintain the original camera viewpoint. Add realistic floodwaters and overcast, stormy weather. The final image should look like a photo taken in the rain, with imperfections like a few water droplets and a smear across a portion of the lens.